ground in 1.90 s. You ay to get to the
Answer #1

The brick is in free-fall and so it is constant acceleration with a = g = 9.8 m/s2.

Since it is falling downwards, its vertical component of velocity will be negative and it will increase linearly with time.

Therefore, vy vs t graph will be a straight line with positive slope m = 9.8 m/s2 as long as v is expressed as negative (which is the case in the above graph).

y vs t graph will be a parabola since its slope dy/dt = vy itself is changing with time.

Time taken to reach the ground = 1.9s

so, the height from which the brick is dropped is: h = ut + (1/2)gt2 = 0 + (1/2)(9.8)(1.9)2 = 17.68 m.